with a tailwind, a bird flew at a ground speed of 3 miles per hour. flying the same path against the same wind, the bird travels at a ground speed of 1.5 miles per hour. what is the bird's air speed? what is the wind speed?
it would be greatly appreciated if you showed the work.
thanks for all your help

Answer by stanbon(75887)

with a tailwind, a bird flew at a ground speed of 3 miles per hour. flying the same path against the same wind, the bird travels at a ground speed of 1.5 miles per hour. what is the bird's air speed? what is the wind speed?
------
Equations:
b + w = 3 mph
b - w = 1.5 mph
----
Add and solve for "b":
2b = 4.5
b = 2.25 mph (bird's speed in still air)
-----
Solve for "w":
b + w = 3 mph
2.25 + w = 3 mph
w = 0.75 mph
===================
Cheers,
Stan H.
===================